OpenStudy (anonymous):

How would you solve this equation 5.9m + 6.8 = 17.36 - 13.3m ??¿

OpenStudy (anonymous):

"you" I think it will be you to solve this..

OpenStudy (anonymous):

I started with combing the like terms

OpenStudy (anonymous):

Good start.. Great.. So, show what you got as your first step??

OpenStudy (anonymous):

That 5.9m added to - 13.3 got me -7.4

OpenStudy (anonymous):

-7.4m - 6.8 = 17.36 -17.36 -17.36 -7.4m -10.56 _____ _____ -7.4 -7.4 = 1.427 m = 1.427

OpenStudy (anonymous):

But that doesnt look right to me, like i missed a step. Or started off wrong

OpenStudy (anonymous):